40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 3 23:42:34 2010, Jenne, Update, PSL, PSL update of the day 
    Reply  Mon Sep 6 01:46:58 2010, rana, Update, PSL, PSL update of the day 
Message ID: 3525     Entry time: Mon Sep 6 01:46:58 2010     In reply to: 3523
Author: rana 
Type: Update 
Category: PSL 
Subject: PSL update of the day 

To bypass the polarizer issue, I just used cubes. One I took from the FSS-Refcav path and the other from the power control part of the old MOPA, just downstream of the MOPA's periscope.

We'll swab these out with the thin-film polarizers after we get the mounts made.

With the cubes in, I also installed the Faraday + its 1/2-wave plate. The transmission looks good and we're getting into the PMC and its flashing a TEM00 mode sometimes. I set up a signal generator to drive the SLOW actuator by 1 FSR at 0.1 Hz.

I have set up a PMC transmission camera and transPD so that its easy to align. The flashing mode already allows us to align most of the rest of the table (except FSS).

Our next step should be to run the cables for locking the PMC:

  1. RF cables to the PMC_REFL
  2. Dsub for the PMC RFPD
  3. HV cable between the PMC servo board and the PMC PZT (why is this not RED? we have to make sure to abide by the cabling color code).
  4. RF cable from 35.5 MHz Frequency Reference card to the PMC EOM via the RF summing box on the table.

On Tuesday, we need to make sure that all of our mounts' drawings are in the cue for the shop. I'll put the list of mounts onto the PSL upgrade wiki page.

We also have to come up with a plan for wiring some of the 2W NPRO's channels into the cross-connect so that we can have some laser channels recorded by EPICS.

ELOG V3.1.3-